# Even Faster Algorithm for the Chamfer Distance

### v1

@@ 1 Introduction
[…]
@@ 1.1 Our techniques
In the second step, the algorithm estimates the Chamfer distance via importance sampling. Specifically, the algorithm samples $T$ points from $A$, such that the probability of sampling $a$ is proportional to the estimate $\mathcal{D}_{a}$. For each sampled point $a$, the distance $\mathsf{opt}_{a}$ from $a$ to its nearest neighbor in $B$ is computed directly in $\mathcal{O}(nd)$ time. The final estimate of the Chamfer distance is equal to the weighted average the $T$ values $\mathsf{opt}_{a}$ It can be shown that if the number of samples $T$ is equal to the distortion $\mathcal{O}(\log n)$ of the estimates $\mathcal{D}_{a}$, this yields a constant factor approximation to the Chamfer distance from $A$ to $B$. The overall cost of the second step is $\mathcal{O}(Tnd)=\mathcal{O}(nd\log n)$, i.e., asymptotically the same as the cost of the first step.
@@ 2 Preliminaries
In this paper, we consider the regime where the approximation factor $\varepsilon\geq\frac{\log^{2}n}{\sqrt{n}}$. Note that otherwise, an $\mathcal{O}(nd/\varepsilon^{2})$ time bound would be close to the runtime of a naive exact computation.
@@ 3 Quadtree
[…]
frametitle= QuadTree , skipabove=, skipbelow=, frametitleaboveskip=-0.7frametitlealignment= {mdframed}
Input: Two size-$n$ subsets $Q:=\{q_{i}\}_{i\in[n]}$ and $P:=\{p_{i}\}_{i\in[n]}$ of a metric space $(\mathbb{R}^{d},\lVert\cdot\rVert_{1})$, such that $Q,P\subset[0,\alpha]^{d}$ for some bound $\alpha={\rm poly}(n)$.
Output: A set of $n$ values $\{\mathcal{D}_{i}\}_{i\in[n]}$, such that every $\mathcal{D}_{i}\in\mathbb{R}$ satisfies $\mathcal{D}_{i}\geq\mathsf{opt}^{P}_{q_{i}}$.
  1. 1.
    Let $t=\lceil\log(\alpha)\rceil+1$. Sample two uniformly random points $z,z^{\prime}\sim[0,2^{t-1}]^{d}$. For any point $x\in[0,\alpha]^{d}$, define
    $h(x):=(\lceil\vec{x}_{1}+\vec{z}_{1}\rceil,\lceil\vec{x}_{2}+\vec{z}_{2}\rceil,\cdots,\lceil\vec{x}_{d}+\vec{z}_{d}\rceil),$
    $h^{\prime}(x):=(\lceil\vec{x}_{1}+\vec{z^{\prime}}_{1}\rceil,\lceil\vec{x}_{2}+\vec{z^{\prime}}_{2}\rceil,\cdots,\lceil\vec{x}_{d}+\vec{z^{\prime}}_{d}\rceil),$
    where $\vec{x}_{i},\vec{z}_{i},\vec{z^{\prime}}_{i}$ are the $i$-th coordinates of $x,z,z^{\prime}$, respectively.
  2. 2.
    For each $x\in Q\cup P$:
    • Compute $h(x)$ and write each element of $h(x)$ as a $t$-bit binary string. Then $h(x)$ can be viewed as a $d$-by-$t$ binary matrix stored in the row-major order, whose $(i,j)$-th entry is the $j$-th significant bit of the $i$-th element of $h(x)$. Transpose this matrix and concatenate the rows of the transpose. Denote the resulting binary string as $h(x)^{\top}$.
    • Similarly, compute $h^{\prime}(x)^{\top}$.
  3. 3.
    Use $h(x)^{\top}$ as keys to sort all $x\in Q\cup P$. Also, use $h^{\prime}(x)^{\top}$ as keys to sort all $x\in Q\cup P$.
  4. 4.
    For each $q_{i}\in Q$:
    • Use the sort to find a $p\in P$ that maximizes the length $l$ of the longest common prefix of $h(q_{i})^{\top}$ and $h(p)^{\top}$. Similarly, find a $p^{\prime}\in P$ that maximizes the length $l^{\prime}$ of the longest common prefix of $h^{\prime}(q_{i})^{\top}$ and $h^{\prime}(p^{\prime})^{\top}$.
    • If $l\geq l^{\prime}$ then output $\mathcal{D}_{i}:=\lVert q_{i}-p\rVert_{1}$; otherwise, output $\mathcal{D}_{i}:=\lVert q_{i}-p^{\prime}\rVert_{1}$.
@@
Correctness: For any $x\in[0,\alpha]^{d}$ and any integer $k$ such that $0\leq k\leq t$, let $,h_{k}(x):=(\lceil\frac{\vec{x}_{1}+\vec{z}_{1}}{2^{k}}\rceil,\lceil\frac{\vec{x}_{2}+\vec{z}_{2}}{2^{k}}\rceil,$$\cdots,\lceil\frac{\vec{x}_{d}+\vec{z}_{d}}{2^{k}}\rceil)$, where $z$ is the random point drawn on Line 1 in Figure 1. Observe that $h_{k}(x)$ is related to the prefix of $h(x)^{\top}$.
@@
Let $q,p\in[0,\alpha]^{d}$ be arbitrary. For any integer $k$ such that $0\leq k\leq t$, $h_{k}(q)=h_{k}(p)$ if and only if $h(q)^{\top}$ and $h(p)^{\top}$ share a common prefix of length at least $d(t-k)$.
@@
If $h(q)^{\top}$ and $h(p)^{\top}$ share a common prefix of length at least $d(t-k)$, then in hashes $h(q)$ and $h(p)$, the first $(t-k)$ bits of all $d$ coordinates are the same. $h_{k}(q)$ and $h_{k}(p)$ compute exactly these bits, thus $h_{k}(q)=h_{k}(p)$. The reverse direction holds symmetrically. ∎
@@
Figure 2: A pictorial example for $8$-by-$8$ square matrix and $w=4$. The left figure (a) shows how the Transpose algorithm handles the rows of $\tilde{M}$ in lines 2a and 3a. The right figure (b) illustrates the transpose outcome.
[…]
frametitle= Transpose , skipabove=, skipbelow=, frametitleaboveskip=-0.7frametitlealignment= {mdframed}
Input: An $I$-by-$J$ bit matrix $M$ where $I\leq J$ are powers of $2$. An integer $w$ that is a power of $2$ and $2\leq w\leq I$.
Output: An $I$-by-$J$ matrix $M^{\prime}$ such that if it is partitioned into $w$-by-$w$ square submatrices, then each submatrix is the transpose of the corresponding submatrix of $M$ at the same coordinates.
  1. 1.
    Let
    $\tilde{M}=\begin{cases}M&\text{if $w$=2}\\ \texttt{Transpose}(M,w/2)&\text{otherwise}\end{cases}$
    be zero-indexed and $\tilde{M}[i,j]$ is its $(i,j)$-th entry.
  2. 2.
    For each integer $i$ such that $0\leq i<I$:
    1. (a)
      Compute a $J$-bit binary string $b_{i}$ such that for $j:0\leq j<J$, its $j$-th bit $b_{i}[j]=\begin{cases}\tilde{M}[i,j]&\text{if $(j\mod w)<w/2$}\\ 0&\text{otherwise}\end{cases}$.
      Also, compute a string $\overline{b_{i}}[j]=\begin{cases}0&\text{if $(j\mod w)<w/2$}\\ \tilde{M}[i,j]&\text{otherwise}\end{cases}$.
  3. 3.
    Define an $I$-by-$J$ matrix $M^{\prime}$, such that for each integer $0\leq i<I$:
    1. (a)
      Let the $i$-th row of $M^{\prime}$ be $\begin{cases}b_{i}+b_{i+(w/2)}\gg(w/2)&\text{if $(i\mod w)<w/2$}\\ \overline{b_{i}}+\overline{b_{i-(w/2)}}\ll(w/2)&\text{if $(i\mod w)\geq w/2$}\end{cases}$,
      where $\gg(w/2)$ (resp. $\ll$) denote the operation of shifting a string to the right (resp. left) by $w/2$ bits.
  4. 4.
    Output $M^{\prime}$.
@@
Figure 3: A pictorial example that shows the behavior of Line 2a and 3a of the Concatenate algorithm on a $4$-by-$4w$ matrix.
[…]
frametitle= Concatenate , skipabove=, skipbelow=, frametitleaboveskip=-0.7frametitlealignment= {mdframed}
Input: An $I$-by-$J$ bit matrix $M$ where $I\leq J$ are powers of $2$. An integer $w$ that is a power of $2$ and $I\leq w\leq J$.
Output: An $IJ$-bit string $B$ such that if it is partitioned into $w$-bit blocks, then the $u$-th block (zero-indexed from left to right) are bits on the $(u\mod I)$-th row of $M$ from column $w\cdot\lfloor u/I\rfloor$ to $w\cdot\lfloor u/I\rfloor+w$.
  1. 1.
    If $w=I$ then output $B=M$.
  2. 2.
    For each integer $i$ such that $0\leq i<I$:
    1. (a)
      Partition the $i$-th row of $M$ into $w$-bit blocks, denoted as $\overbrace{[\underbrace{b_{i,1}}_{w}\mid b_{i,2}\mid\ldots\mid b_{i,J/w}]}^{J}$. Compute a $2J$-bit string $b_{i}=[b_{i,1}\mid 0_{w}\mid b_{i,2}\mid 0_{w}\mid\ldots\mid b_{i,J/w}\mid 0_{w}]$, where $0_{w}$ is a $w$-bit all-zero string.
  3. 3.
    Define an $I/2$-by-$2J$ matrix $M^{\prime}$, such that for each integer $20\leq i<I/$ 2:
    1. (a)
      Let the $i$-th row of $M^{\prime}$ be $b_{2i}+b_{2i+1}\gg w$, where $\gg w$ is the operation of shifting a string to the right by $w$ bits.
  4. 4.
    Output $\texttt{Concatenate}(M^{\prime},2w)$.
The correctness of the Concatenate algorithm can again be observed by inducting on the logarithm of $w$. Line 2a and 3a can be done using $\mathcal{O}((J/w)\cdot\lceil w/\log n\rceil)$ operations on words, and both lines are repeated for $I$ times in each recursive call. Therefore, the total runtime is
@@ 4 Tournament
In this section, we compute the $2$-approximation of the nearest neighbor distances for logarithmically many queries. We do so using a depth-$2$ tournament. In one branch of the tournament tree, we sample a small set $\overline{S}$ of input points uniformly at random. In the other branch, we partition input points into random groups, project them to a lower dimensional space, and then collect the nearest neighbor in the projected space in every group as a set $\tilde{S}$. The final output of the tournament is the nearest neighbor among points in $\overline{S},\tilde{S}$ in the original space. Intuitively, if there are many ($2$-approximate) near neighbors, then a random subset $\overline{S}$ should contain one of them. And when there are few, these neighbors are likely to be assigned to different random groups, in which case the true nearest neighbor should be collected to $\tilde{S}$.
@@
Let $x,y\in\mathbb{R}^{d}$ and $20<c<1/$ 2. Sample $r$ random vectors $v_{1},v_{2},\cdots,v_{r}\sim(\mathsf{Cauchy}(0,1))^{d}$. With probability at least $1-2e^{-rc^{2}/50}$, $\mathsf{med}\{|v_{i}\cdot(x-y)|:i\in[r]\}\in(1\pm c)\lVert x-y\rVert_{1}$.
@@
[…]
frametitle= Tournament , skipabove=, skipbelow=, frametitleaboveskip=-0.7frametitlealignment= {mdframed}
Input: A set of $t$ queries $\{q_{i}\}_{i\in[t]}$ and a set of $n$ points $P$, which are both subsets of a metric space $(\mathbb{R}^{d},\|\cdot\|_{1})$.
Output: A set of $t$ values $\{\mathcal{D}_{i}\}_{i\in[t]}$, such that every $\mathcal{D}_{i}\in\mathbb{R}$ satisfies $\mathcal{D}_{i}\geq\mathsf{opt}^{P}_{q_{i}}$.
@@
  1. 1.
    Let $r\geq 800(2\log t+\log\log n)$.
  2. 2.
    For each $j\in[r]$, draw $v_{j}\sim(\mathsf{Cauchy}(0,1))^{d}$, compute $v_{j}\cdot p$ for all points $p\in P$, and store all $v_{j}$ and $v_{j}\cdot p$.
  3. 3.
    Randomly partition $P$ into $n/\log n$ subsets $P_{1},P_{2},\cdots,P_{n/\log n}$, each of size $\log n$.
@@
For each query $q:=q_{i}$ for $i\in[t]$:
  1. 1.
    Compute $v_{j}\cdot q$ for all $j\in[r]$.
  2. 2.
    Let $\tilde{S}$ be an empty set. For each $k\in[n/\log n]$:
    • Compute $\mathsf{med}_{p}:=\mathsf{med}\{|v_{j}\cdot(q-p)|:j\in[r]\}$ for every $p\in P_{k}$.
    • Find $p:=\operatorname*{arg\,min}_{p\in P_{k}}\{\mathsf{med}_{p}\}$ and add it into $\tilde{S}$.
  3. 3.
    Find $\tilde{p}=\operatorname*{arg\,min}_{p\in\tilde{S}}\lVert q-p\rVert_{1}$ by computing and comparing all exact distances $\lVert q-p\rVert_{1}$ for $p\in\tilde{S}$.
  4. 4.
    Let $\overline{S}$ be a set of $w\geq 90t\log t\log n$ samples drawn uniformly at random from $P$. Find a point $\overline{p}=\operatorname*{arg\,min}_{p\in\overline{S}}\lVert q-p\rVert_{1}$ by computing and comparing all exact distances $\lVert q-p\rVert_{1}$ for $p\in\overline{S}$.
  5. 5.
    Output $\mathcal{D}_{i}:=\min{(\lVert q-\tilde{p}\rVert_{1},\lVert q-\overline{p}\rVert_{1})}$.
@@
Fix a query $q:=q_{i}$. Let $S$ denote the set of all $2$-approximate nearest neighbors to $q$, i.e., $S:=\{p\in P:\lVert q-p\rVert_{1}\leq 2\mathsf{opt}_{q}\}.$ We prove the correctness of the algorithm by casing on the size of $S$.
Lemma 4.2 (Case 1).
If $|S|\geq\frac{n}{30t\log n}$ then with probability at least $1-\frac{1}{10t}$ we have $\overline{p}\in S$.
Proof.
With probability $\frac{1}{30t\log n}$, a random sample from $P$ is in $S$. Therefore, for $\overline{S}$ containing $w$ independent samples, at least one of them is in $S$ with probability $1-(1-\frac{1}{30t\log n})^{w}\geq 1-e^{-w/(30t\log n)}$. Moreover, when this happens, it must be that $\overline{p}\in S$. Setting $w\geq 90t\log t\log n$ gives the desired probability.
Lemma 4.3 (Case 2).
If $|S|<\frac{n}{30t\log n}$ then with probability at least $1-\frac{1}{10t}$, $\tilde{p}\in S$.
Let $p^{*}\in P$ denote a nearest neighbor of $q$, i.e. $\lVert q-p^{*}\rVert_{1}=\mathsf{opt}_{q}$. To prove Lemma 4.3, we first make the following observation:
@@
In Line 3 of the data structure building procedure, the point $p^{*}$ is assigned to one of the subsets $,P^{*}\in\{P_{1},P_{2},$$\cdots,P_{n/\log n}\}$. If $|S|<\frac{n}{30t\log n}$, then one can show that $p^{*}$ is likely the only $2$-approximate nearest neighbor in $P^{*}$. Conditioned on this, we can use Lemma 4.4 to show that $p^{*}$ is added into $\tilde{S}$ with high probability.
Lemma 4.5.
If $|S|<\frac{n}{30t\log n}$, then with probability at least $1-\frac{1}{10t}$, $p^{*}$ is added into $\tilde{S}$ on Line 2 when processing the query $q$.
Proof.
The set $P^{*}$ contains $\log n$ points. Since we randomly partition $P$ into $P_{1},\cdots,P_{n/\log n}$, $P^{*}\setminus\{p^{*}\}$ is a uniformly random subset of $P$. When $|S|<\frac{n}{30t\log n}$, $\text{\bf Pr}\bigl{[}(P^{*}\setminus\{p^{*}\})\cap S=\varnothing\bigr{]}\geq 1-\frac{1}{20t}$. Conditioned on this event, we have
$\text{\bf Pr}\bigl{[}\exists p\in(P^{*}\setminus\{p^{*}\}):\mathsf{med}_{p}\leq\mathsf{med}_{p^{*}}\bigr{]}\leq\frac{2|P^{*}|}{t^{2}\log n}\leq\frac{1}{20t}$
by Lemma 4.4, as long as $t\geq 40$. Therefore, with probability at least $1-\frac{1}{10t}$, $p^{*}$ has the smallest median of projected distance to $q$, and thus must be added to $\tilde{S}$.
Lemma 4.3 is a direct corollary of Lemma 4.5.
@@
$p^{*}\in\tilde{S}$ implies that if we compute $\tilde{p}=\operatorname*{arg\,min}_{p\in\tilde{S}}\lVert q-p\rVert_{1}$, we are guaranteed to find a nearest neighbor of $q$, which is clearly in $S$.
Combining Lemma 4.2 and 4.3 we get the following correctness guarantee:
@@
For preprocessing, the algorithm projects all points in $P$ using $r$ projections, which takes $\mathcal{O}(n\cdot d\cdot r)$ time. To process a query $q$, we first take $\mathcal{O}(dr)$ time to project $q$. We then count the number of comparisons we make to find the minimums of medians, which is $\mathcal{O}((n/\log n)\log n\cdot r)$ using a linear-time median selection algorithm [BFP+73]. Each comparison can be done in $\mathcal{O}(1)$ time given that $v_{j}\cdot p$ and $v_{j}\cdot q$ for all $j\in[r]$ and $p\in P$ are stored. Finally, we do a linear scan over $\tilde{S}$ and $\overline{S}$, which takes $\mathcal{O}(d(\log n+w))$ time.
We plug in $r=\mathcal{O}(\log t+\log\log n)$ and $w=\mathcal{O}(t\log t\log n)$. For $t$ queries, the total runtime is $\mathcal{O}(n(d+t)(\log t+\log\log n)+dt^{2}\log t\log n)$. ∎
For our purpose of estimating the Chamfer distance, we will apply the Tournament algorithm with a number of queries $t=\Theta({{D}}/\varepsilon^{2})$ for ${{D}}=\min{(d,3\log n)}$ and some $\varepsilon>0$ satisfying $\varepsilon^{-2}=\mathcal{O}(\frac{n}{\log^{4}n})$. Under this setting, the runtime is dominated by the first additive term of Theorem 4.7, which is at most $\mathcal{O}(nd(\log\log n+\log\frac{1}{\varepsilon})/\varepsilon^{2})$.

### v2

@@ 1 Introduction
where $\|\cdot\|$ is the underlying norm defining the distance between the points. Chamfer distance and its variant, the Relaxed Earth Mover Distance [KSKW15, AM19], are widely used metrics for quantifying the distance between two sets of points. These measures are especially popular in fields such as machine learning (e.g.,[KSKW15, WCL+19]) and computer vision (e.g.,[AS03, SMFW04, FSG17, JSQJ18]). A closely related notion of “the sum of maximum similarities”, where $\min_{b\in B}\|a-b\|$ is replaced by $\max_{b\in B}a\cdot b$, has been recently popularized by the ColBERT system [KZ20]. Efficient subroutines for computing Chamfer distances are provided in prominent libraries including Pytorch [pyt23], PDAL [pda23] and Tensorflow [ten23]. In many applications (e.g., see  [KSKW15]), Chamfer distance is favored as a faster alternative to the more computationally intensive Earth-Mover Distance or Wasserstein Distance.
@@ 1.1 Our techniques
In the second step, the algorithm estimates the Chamfer distance via importance sampling. Specifically, the algorithm samples $T$ points from $A$, such that the probability of sampling $a$ is proportional to the estimate $\mathcal{D}_{a}$. For each sampled point $a$, the distance $\mathsf{opt}_{a}$ from $a$ to its nearest neighbor in $B$ is computed directly in $\mathcal{O}(nd)$ time. The final estimate of the Chamfer distance is equal to the weighted average the $T$ values $\mathsf{opt}_{a}$ . It can be shown that if the number of samples $T$ is equal to the distortion $\mathcal{O}(\log n)$ of the estimates $\mathcal{D}_{a}$, this yields a constant factor approximation to the Chamfer distance from $A$ to $B$. The overall cost of the second step is $\mathcal{O}(Tnd)=\mathcal{O}(nd\log n)$, i.e., asymptotically the same as the cost of the first step.
@@ 2 Preliminaries
In this paper, we consider the regime where the approximation factor $\varepsilon\geq\sqrt{\frac{\log n}{n}}$. Note that otherwise, an $\mathcal{O}(nd/\varepsilon^{2})$ time bound would be close to the runtime of a naive exact computation.
@@ 3 Quadtree
Input: Two size-$n$ subsets $Q:=\{q_{i}\}_{i\in[n]}$ and $P:=\{p_{i}\}_{i\in[n]}$ of a metric space $(\mathbb{R}^{d},\lVert\cdot\rVert_{1})$, such that $Q,P\subset[0,\alpha]^{d}$ for some bound $\alpha={\rm poly}(n)$. Output: A set of $n$ values $\{\mathcal{D}_{i}\}_{i\in[n]}$, such that every $\mathcal{D}_{i}\in\mathbb{R}$ satisfies $\mathcal{D}_{i}\geq\mathsf{opt}^{P}_{q_{i}}$. 1. Let $t=\lceil\log(\alpha)\rceil+1$. Sample two uniformly random points $z,z^{\prime}\sim[0,2^{t-1}]^{d}$. For any point $x\in[0,\alpha]^{d}$, define $h(x):=(\lceil\vec{x}_{1}+\vec{z}_{1}\rceil,\lceil\vec{x}_{2}+\vec{z}_{2}\rceil,\cdots,\lceil\vec{x}_{d}+\vec{z}_{d}\rceil),$ $h^{\prime}(x):=(\lceil\vec{x}_{1}+\vec{z^{\prime}}_{1}\rceil,\lceil\vec{x}_{2}+\vec{z^{\prime}}_{2}\rceil,\cdots,\lceil\vec{x}_{d}+\vec{z^{\prime}}_{d}\rceil),$ where $\vec{x}_{i},\vec{z}_{i},\vec{z^{\prime}}_{i}$ are the $i$-th coordinates of $x,z,z^{\prime}$, respectively. 2. For each $x\in Q\cup P$: Compute $h(x)$ and write each element of $h(x)$ as a $t$-bit binary string. Then $h(x)$ can be viewed as a $d$-by-$t$ binary matrix stored in the row-major order, whose $(i,j)$-th entry is the $j$-th significant bit of the $i$-th element of $h(x)$. Transpose this matrix and concatenate the rows of the transpose. Denote the resulting binary string as $h(x)^{\top}$. Similarly, compute $h^{\prime}(x)^{\top}$. 3. Use $h(x)^{\top}$ as keys to sort all $x\in Q\cup P$. Also, use $h^{\prime}(x)^{\top}$ as keys to sort all $x\in Q\cup P$. 4. For each $q_{i}\in Q$: Use the sort to find a $p\in P$ that maximizes the length $l$ of the longest common prefix of $h(q_{i})^{\top}$ and $h(p)^{\top}$. Similarly, find a $p^{\prime}\in P$ that maximizes the length $l^{\prime}$ of the longest common prefix of $h^{\prime}(q_{i})^{\top}$ and $h^{\prime}(p^{\prime})^{\top}$. If $l\geq l^{\prime}$ then output $\mathcal{D}_{i}:=\lVert q_{i}-p\rVert_{1}$; otherwise, output $\mathcal{D}_{i}:=\lVert q_{i}-p^{\prime}\rVert_{1}$.
@@
Correctness: For any $x\in[0,\alpha]^{d}$ and any integer $k$ such that $0\leq k\leq t$, let $h_{k}(x):=(\lceil\frac{\vec{x}_{1}+\vec{z}_{1}}{2^{k}}\rceil,\lceil\frac{\vec{x}_{2}+\vec{z}_{2}}{2^{k}}\rceil,$ $\cdots,\lceil\frac{\vec{x}_{d}+\vec{z}_{d}}{2^{k}}\rceil)$, where $z$ is the random point drawn on Line 1 in Figure 1. Observe that $h_{k}(x)$ is related to the prefix of $h(x)^{\top}$.
@@
Let $q,p\in[0,\alpha]^{d}$ be arbitrary. For any integer $k$ such that $0\leq k\leq t$, $h_{k}(q)=h_{k}(p)$ if and only if $h(q)^{\top}$ and $h(p)^{\top}$ share a common prefix of length at least d(tk)d(t-k).
@@
If $h(q)^{\top}$ and $h(p)^{\top}$ share a common prefix of length at least d(tk)d(t-k), then in hashes $h(q)$ and $h(p)$, the first (tk)(t-k) bits of all $d$ coordinates are the same. $h_{k}(q)$ and $h_{k}(p)$ compute exactly these bits, thus $h_{k}(q)=h_{k}(p)$. The reverse direction holds symmetrically. ∎
@@
Figure 2: A pictorial example for 88-by-88 square matrix and $w=4$. The left figure (a) shows how the Transpose algorithm handles the rows of $\tilde{M}$ in lines 2a and 3a. The right figure (b) illustrates the transpose outcome.
Input: An II-by-$J$ bit matrix $M$ where $I\leq J$ are powers of $2$. An integer $w$ that is a power of $2$ and $2\leq w\leq I$. Output: An II-by-$J$ matrix $M^{\prime}$ such that if it is partitioned into $w$-by-$w$ square submatrices, then each submatrix is the transpose of the corresponding submatrix of $M$ at the same coordinates. 1. Let $\tilde{M}=\begin{cases}M&\text{if $w$=2}\\ \texttt{Transpose}(M,w/2)&\text{otherwise}\end{cases}$ be zero-indexed and $\tilde{M}[i,j]$ is its $(i,j)$-th entry. 2. For each integer $i$ such that $0\leq i<I$: (a) Compute a $J$-bit binary string $b_{i}$ such that for $j:0\leq j<J$, its $j$-th bit $b_{i}[j]=\begin{cases}\tilde{M}[i,j]&\text{if $(j\mod w)<w/2$}\\ 0&\text{otherwise}\end{cases}$. Also, compute a string $\overline{b_{i}}[j]=\begin{cases}0&\text{if $(j\mod w)<w/2$}\\ \tilde{M}[i,j]&\text{otherwise}\end{cases}$. 3. Define an II-by-$J$ matrix $M^{\prime}$, such that for each integer $0\leq i<I$: (a) Let the $i$-th row of $M^{\prime}$ be $\begin{cases}b_{i}+b_{i+(w/2)}\gg(w/2)&\text{if $(i\mod w)<w/2$}\\ \overline{b_{i}}+\overline{b_{i-(w/2)}}\ll(w/2)&\text{if $(i\mod w)\geq w/2$}\end{cases}$, where $\gg(w/2)$ (resp. $\ll$) denote the operation of shifting a string to the right (resp. left) by $w/2$ bits. 4. Output $M^{\prime}$.
@@
Figure 3: A pictorial example that shows the behavior of Line 2a and 3a of the Concatenate algorithm on a 44-by-$4w$ matrix.
Input: An II-by-$J$ bit matrix $M$ where $I\leq J$ are powers of $2$. An integer $w$ that is a power of $2$ and $I\leq w\leq J$. Output: An IJIJ-bit string $B$ such that if it is partitioned into $w$-bit blocks, then the $u$-th block (zero-indexed from left to right) are bits on the $(u\mod I)$-th row of $M$ from column $w\cdot\lfloor u/I\rfloor$ to $w\cdot\lfloor u/I\rfloor+w$. 1. If $w=I$ then output $B=M$. 2. For each integer $i$ such that $0\leq i<I$: (a) Partition the $i$-th row of $M$ into $w$-bit blocks, denoted as $\overbrace{[\underbrace{b_{i,1}}_{w}\mid b_{i,2}\mid\ldots\mid b_{i,J/w}]}^{J}$. Compute a $2J$-bit string $b_{i}=[b_{i,1}\mid 0_{w}\mid b_{i,2}\mid 0_{w}\mid\ldots\mid b_{i,J/w}\mid 0_{w}]$, where $0_{w}$ is a $w$-bit all-zero string. 3. Define an $I/2$-by-$2J$ matrix $M^{\prime}$, such that for each integer $0\leq i<I/2$: (a) Let the $i$-th row of $M^{\prime}$ be $b_{2i}+b_{2i+1}\gg w$, where $\gg w$ is the operation of shifting a string to the right by $w$ bits. 4. Output $\texttt{Concatenate}(M^{\prime},2w)$.
The correctness of the Concatenate algorithm can again be observed by inducting on the logarithm of $w$. Line 2a and 3a can be done using $\mathcal{O}((J/w)\cdot\lceil w/\log n\rceil)$ operations on words, and both lines are repeated for II times in each recursive call. Therefore, the total runtime is
@@ 4 Tournament
In this section, we compute the $2$-approximation of the nearest neighbor distances for logarithmically many queries. We do so using a depth-$2$ tournament: we first partition input points into random groups, project them to a lower dimensional space, and collect the nearest neighbor in the projected space in every group as a set $\tilde{S}$. Then the final output of the tournament is the nearest neighbor among points in $\tilde{S}$ in the original space. Intuitively, because each random group is small, the true nearest neighbor could only lose to another near neighbor in the first step. Then in the second step, $\tilde{S}$ should contain at least one near neighbor.
@@
Let $x,y\in\mathbb{R}^{d}$ and $0<c<1/2$. Sample $r$ random vectors $v_{1},v_{2},\cdots,v_{r}\sim(\mathsf{Cauchy}(0,1))^{d}$. With probability at least $1-2e^{-rc^{2}/50}$, $\mathsf{med}\{|v_{i}\cdot(x-y)|:i\in[r]\}\in(1\pm c)\lVert x-y\rVert_{1}$.
@@
Input: A set of $t$ queries $\{q_{i}\}_{i\in[t]}$ and a set of $n$ points $P$, which are both subsets of a metric space $(\mathbb{R}^{d},\|\cdot\|_{1})$. Output: A set of $t$ values $\{\mathcal{D}_{i}\}_{i\in[t]}$, such that every $\mathcal{D}_{i}\in\mathbb{R}$ satisfies $\mathcal{D}_{i}\geq\mathsf{opt}^{P}_{q_{i}}$.
@@
1. Let $r\geq 800(2\log t+\log\log n)$. 2. For each $j\in[r]$, draw $v_{j}\sim(\mathsf{Cauchy}(0,1))^{d}$, compute $v_{j}\cdot p$ for all points $p\in P$, and store all $v_{j}$ and $v_{j}\cdot p$. 3. Randomly partition $P$ into $n/\log n$ subsets $P_{1},P_{2},\cdots,P_{n/\log n}$, each of size $\log n$.
@@
For each query $q:=q_{i}$ for $i\in[t]$: 1. Compute $v_{j}\cdot q$ for all $j\in[r]$. 2. Let $\tilde{S}$ be an empty set. For each $k\in[n/\log n]$: Compute $\mathsf{med}_{p}:=\mathsf{med}\{|v_{j}\cdot(q-p)|:j\in[r]\}$ for every $p\in P_{k}$. Find $p:=\operatorname*{arg\,min}_{p\in P_{k}}\{\mathsf{med}_{p}\}$ and add it into $\tilde{S}$. 3. Output $\mathcal{D}_{i}:=\min_{p\in\tilde{S}}\lVert q-p\rVert_{1}$ by computing and comparing all exact distances $\lVert q-p\rVert_{1}$ for $p\in\tilde{S}$.
@@
We fix a query $q:=q_{i}$.
Lemma 4.2.
With probability at least $1-\frac{1}{10t}$, $\mathcal{D}_{i}\leq 2\mathsf{opt}_{q}$.
We let $S$ denote the set of all $2$-approximate nearest neighbors to $q$, i.e., $S:=\{p\in P:\lVert q-p\rVert_{1}\leq 2\mathsf{opt}_{q}\}$, and let $p^{*}\in P$ denote a nearest neighbor of $q$, i.e. $\lVert q-p^{*}\rVert_{1}=\mathsf{opt}_{q}$. To prove Lemma 4.2, we first make the following observation:
@@
In Line 3 of the data structure building procedure, the point $p^{*}$ is assigned to one of the subsets $P^{*}\in\{P_{1},P_{2},$ $\cdots,P_{n/\log n}\}$. Focusing on this subset $P^{*}$, we can show that with high probability, either $p^{*}$ is added into $\tilde{S}$, or $p^{*}$ loses to another 2-approximate nearest neighbor. In both cases, the data structure is guaranteed to output an 2-approximation.
@@
$P^{*}\setminus S$ contains at most $|P^{*}|=\log n$ points. Applying Lemma 4.3, we get $\mathsf{med}_{p^{*}}\geq\mathsf{med}_{p}$ simultaneously for all $p\in P^{*}\setminus S$ with probability at least $1-\frac{2\log n}{t^{2}\log n}\geq 1-\frac{1}{10t}$ (as long as $t\geq 20$). Conditioned on this, $\operatorname*{arg\,min}_{p\in P^{*}}\{\mathsf{med}_{p}\}$ must be either $p^{*}$ or some other element of $S$. Thus on line 3 of the algorithm, $\tilde{S}$ contains at least one element of $S$, so the final output $\min_{p\in\tilde{S}}\|q-p\|_{1}\leq 2\mathsf{opt}_{q}$.
Applying a union bound on Lemma 4.2, we get the correctness guarantee:
@@
For preprocessing, the algorithm projects all points in $P$ using $r$ projections, which takes $\mathcal{O}(n\cdot d\cdot r)$ time. To process a query $q$, we first take $\mathcal{O}(dr)$ time to project $q$. We then count the number of comparisons we make to find the minimums of medians, which is $\mathcal{O}((n/\log n)\log n\cdot r)$ using a linear-time median selection algorithm [BFP+73]. Each comparison can be done in $\mathcal{O}(1)$ time given that $v_{j}\cdot p$ and $v_{j}\cdot q$ for all $j\in[r]$ and $p\in P$ are stored. Finally, we use $\mathcal{O}(d\log n)$ time to do a linear scan over $\tilde{S}$.
We plug in $r=\mathcal{O}(\log t+\log\log n)$. For $t$ queries, the total runtime is $\mathcal{O}(n(d+t)(\log t+\log\log n)+dt\log n)$. ∎
For our purpose of estimating the Chamfer distance, we will apply the Tournament algorithm with a number of queries $t=\Theta({{D}}/\varepsilon^{2})$ for ${{D}}=\min{(d,3\log n)}$ and some $\varepsilon>0$ satisfying $\varepsilon^{-2}=\mathcal{O}(\frac{n}{\log n})$. Under this setting, the runtime is dominated by the first additive term of Theorem 4.5, which is at most $\mathcal{O}(nd(\log\log n+\log\frac{1}{\varepsilon})/\varepsilon^{2})$.